OpenStudy (anonymous):

The coefficient of variation for a sample of values Y1,...,Yn is CV=S/Y bar defined by where Ybar is the mean and S is the standard deviation. Let denote a random sample of size 8 from a normal distribution with mean 0 and variance. a) Obtain the distribution of 8Ybar/s^2 b)What is the distribution of s^2/8ybar^2?

OpenStudy (kropot72):

The variance value has not been included in the question. Please post value of the variance.

OpenStudy (campbell_st):

looks like standard normal, mean 0 standard deviation of 1

OpenStudy (anonymous):

sorry..typing mistakes..variance is 1

OpenStudy (anonymous):

Show that the series \[\sum_{n=0}^{\infty} 1/(1+n^2 x) \] converges uniformly on \[\left[ a, \infty \right]\] for a >0 but does not converges uniformly on \[\left( 0, \infty \right)\].